What is the equation of a parabola with vertex (0,1) that passes through the point (3,9)?
Answer by scott8148(6628)

vertex form of a quadratic is ___ y = a(x - h)^2 + k ; where (h,k) is the vertex

y = a(x - 0)^2 + 1

9 = a(3)^2 + 1 ___ 8 = 9a ___ 8/9 = a

y = (8/9)x^2 + 1